OpenStudy (anonymous):

Why does 1/25 = 2^-5 ?

OpenStudy (anonymous):

and can you explain this rule to me?

OpenStudy (anonymous):

because 2 is raised to a negative power.

OpenStudy (anonymous):

that doesn't really help ..

OpenStudy (anonymous):

thats one of the exponent rules. read your book or use google!

OpenStudy (anonymous):

a negative exponent is the reciprocal of the positive so x^-2 = 1/(x^2) = 1/5 and 5^-2 = 1/25

OpenStudy (anonymous):

^-----This

OpenStudy (anonymous):

yea it a rule, learn it, live it, love it!

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

You will see it again when learning derivatives!

OpenStudy (anonymous):

just try to have fun with it!

OpenStudy (anonymous):

i can't. math is the death of me :(

OpenStudy (anonymous):

no, it just patterns and rules. Like a language.

OpenStudy (anonymous):

wait! wouldn't it equal 5^ -2?

OpenStudy (anonymous):

5^ -2 = 1/5^2 = 1/25

OpenStudy (anonymous):

is that a yes..

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so if the rule was 1/a^m = a^ -m......a and m can switch basically?

OpenStudy (anonymous):

let me be clear x^-b = 1/(x^b), yes?

OpenStudy (anonymous):

so like 1/25 = 1/5^2 = 5^-2 WHICH IS ALSO 2^ -5 so the "a" and "m" switched

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

but am i right...what i said?

OpenStudy (anonymous):

clear now? Thats the rule, among many :)

OpenStudy (anonymous):

please respond to what i said before...

OpenStudy (anonymous):

about a and m switching ?? lemme look at it.

OpenStudy (anonymous):

can the two values switch like that....like 5^ -2 = 2^ -5

OpenStudy (anonymous):

am i right to say that

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

"1/25 = 1/5^2 = 5^-2 WHICH IS ALSO 2^ -5 " this is not true 2^-5 = 1/( 2^5) = 1/32- I misstyped.

OpenStudy (anonymous):

okay. do u mind giving me a few of these to practice...like quizzing me i guess lol

OpenStudy (anonymous):

sure 3^-3

OpenStudy (anonymous):

ummm oh god

OpenStudy (anonymous):

lets see..

OpenStudy (anonymous):

1/3^3

OpenStudy (anonymous):

ugh im wrong

OpenStudy (anonymous):

so far so good- and what is 1/(3^3)?

OpenStudy (anonymous):

that would be 3^-3

OpenStudy (anonymous):

yes, thats true too but what is 3^3?

OpenStudy (anonymous):

27

OpenStudy (anonymous):

so 1/(3^3) = 1/27 right?

OpenStudy (anonymous):

can u give me one where a and m are different numbers

OpenStudy (anonymous):

2^-3

OpenStudy (anonymous):

okay so this would be 1/2^3

OpenStudy (anonymous):

so 1/8

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

yes! so you seem to understand.

OpenStudy (anonymous):

can u do one with very big numbers? well like somewhere in the 200s-300s?

OpenStudy (anonymous):

sure 222^-300

OpenStudy (anonymous):

uh oh.... umm lemme think

OpenStudy (anonymous):

1/222^300

OpenStudy (anonymous):

okay i think i got this now ! just gotta memorize that rule! :D

OpenStudy (anonymous):

are you gonna be online bc i might have more questions :)

OpenStudy (anonymous):

you might want a calculator at this point, The rule x^-n = 1/(x^n) is true for ALL number

OpenStudy (anonymous):

I will be around until Im not :)

OpenStudy (anonymous):

we are not allowed to use calculators on our quiz :(

OpenStudy (anonymous):

yea so 222^-300 might not be on the quiz :P btw the answer is : 1/80517880668393229147886068496149401306972582290411328168978510994785579115588583551827976204875447471474879862267202705113127441091878902354853813041890222152198455342144439084242078428238138714265508683035791505554760104867426467157715278054403172295031898472730484002817659144331717312161700756934610875579137993642195693033512692240606656314618045823043091894260503678798685779973005364739112564184555335918720320643432850355542239025011335644126445607944039212224942030284232383674702362906793619267527992062005620717473752544451951685766644998888627609736993334783701457851056538472011058453524224265800042617872532780902329250444294802910839828000675334422137717517355426310014485465179583860965376

OpenStudy (anonymous):

study hard.

OpenStudy (anonymous):

oh my goodness. are you like a math teacher or something. you are so good at this stuff

OpenStudy (anonymous):

no, Im a student, like you.

OpenStudy (anonymous):

lol i am a junior in high school and i still do not understand this stuff

OpenStudy (anonymous):

practice, practice, practice.

OpenStudy (anonymous):

ugh. i hate math so much. do you know much about logs

OpenStudy (anonymous):

some what is the deal?

OpenStudy (anonymous):

hold on give me a minute

OpenStudy (anonymous):

you should open another question- there a lot of smart people here.

OpenStudy (anonymous):

okay i will. thank you for all your help! hope to talk to you more sometime on the webbiste! :) - Olivia

OpenStudy (anonymous):

sure, I will keep an eye out for your questions :)